OpenStudy (anonymous):

At the beginning of a walk, Roberto and Juanita are 2.2 miles apart. If they leave at the same time and walk in the same direction, Roberto overtakes Juanita in 11 hours. If they walk towards each other, they meet in 1 hour. What are their speeds? Roberto’s speed is mph Juanita’s speed is mph

OpenStudy (saifoo.khan):

Juanita distance = "x" miles ; Roberto distance = "x+2.2" miles Time for both is 11 hrs. Juanita rate = x/11 mph ; Roberto rate = (x+2.2)/11 mph =========================================================== If they walk towards each other, they meet in 1 hour. Juanita distance = y miles ; Roberto distance = 2.2-y miles Time for both is 1 hr. Juanita rate = y mph ; Roberto rate = (2.2-y) mph --- What are their speeds? ================================ Equations: Juanita rate: x/11 = y Roberto rate: (x+2.2)/11 = 2.2-y ------------------------------------ y = x/11 y = 2.2 - [(x+2.2)/11] --------------------------- x/1l = 2.2 -[(x+2.2)/11] x = 24.2 - (x+2.2) 2x = 22 x = 11 miles ===================== Juanita rate = x/11 = 1 mph Robert rate = 2.2-1 = 1.2 mph ====================================